# Topological Data Analysis in smart manufacturing

### v2

@@ Abstract
journal: a Journal
@@ 3 Topological Data Analysis
We consider TDA to consist of three main methods: Mapper [38], PH [17, 39], and Uniform Manifold Approximation and Projection (UMAP) [40]. PH is the prevalent method in TDA, Mapper a topologically guided clustering algorithm, and UMAP a topology-driven dimensionality reduction method. Common to all of them is that (i) the data at hand is converted into a suitable (geo)metric representation and (ii) topological properties are analyzed to extract (novel) information from the given representation. In a remote analogy, where the Fourier analysis tells us something about spectral properties of a signal, TDA extracts information about the shape of various kind of data (e.g., loops, tunnels, cavities, components, et cetera).
@@
During the stage of data preprocessing, data usually is manipulated by a set of parameters. The three aforementioned methods deal with parameters – i.e., the step from input data to a combinatorial object – in different ways. First, Mapper assembles parameters (and their values) in different groups and then clusters the inputs accordingly. This grouping can find latent relations in the dataset. Similarly, UMAP is a dimensionality reduction method. By projecting the data into lower-dimensional ambient space, redundant or not relevant information in the data is compressed. This enables the input to be analyzed more easily. On the other hand, PH overcomes the need of choosing parameter values by considering all possible values at the same time. Instead of outputting information about the shapes in the data for a single parameter value, PH keeps track of how the shapes evolve depending on the variation of parameter(s). PH is especially suitable for automated production since it removes the need to cherrypick a threshold (see for example [43]).
@@ 3.1 Mapper
    an interval cover, cover for brevity, of the image of $f$ (i.e., a collection of possibly overlapping intervals whose union covers the image of $f$), and
@@
We now describe the steps of Mapper (illustrated in Figure 5): first, the input data $X$ is reduced in dimension using the filter function $f$. Then one chooses a cover of the image of $f$. Once the image of $f$ is covered, one takes the preimage of each interval in the cover, i.e., the collection of all data $x\in X$ such that $f(x)\in\mathcal{I}$ for an interval $\mathcal{I}$ in the cover. In other terms, all data points that are mapped to the same interval in the cover are grouped together in the domain. Note that, since the intervals in the cover may overlap (and they usually do), a point $x\in X$ can belong to the preimages of several intervals. Each of these preimages is clustered according to the chosen clustering algorithm. Finally, each of these clusters is represented as a vertex, and two vertices have an edge between them if there is at least one $x\in X$ that belongs both to corresponding clusters.
@@
In practice, the major obstacles to the use of Mapper are the choices of the filter function $f$, cover, and clustering algorithm: the interpretability of the outcome entirely depends on these choices. While a few standard choices are known, a domain expert is usually required to get meaningful insights out of the Mapper pipeline. Nevertheless, Mapper is a powerful, versatile method that can reveal covert connectivity in datasets.
@@ 3.2 Uniform Manifold Approximation and Projection (UMAP)
The work by Hsu et al. [52] exemplifies the application of UMAP. They analyze failures in System Level Tests (SLTs) on integrated circuits. The need for robust interpretative methodologies is driven by the increasing transistor density and complexity in circuit designs, where the high-dimensional nature of the associated data poses a significant challenge. To address this, Hsu et al. [52] project high-dimensional data into a lower-dimensional space using UMAP. The obtained lower-dimensional representation is analyzed with a clustering algorithm, isolating clusters of systematically failing SLTs. These clusters are then leveraged to construct a decision tree, enabling precise identification of failures in individual components. The clustering algorithm is very efficient but sentitive to the curse of dimensionality. Therefore, applying it directly to the data would lower the performances.
@@ 3.3 Persistent Homology (PH)
There are several methods to turn $X_{r}$ into a simplicial complex. We will not elaborate on details, but this choice may affect the interpretation and usually impacts the runtime considerably. Therefore, it is important to choose the more appropriate method for each application to ensure efficient computation [53].
@@
Now that we have a suitable combinatorial object, we need to extract information from it and we do so using homology. Homology is a fundamental concept from algebraic topology, allowing us to identify shapes that cannot be continuously deformed into each other. An extensive treatment is beyond our scope (see for example [17] for more details). Informally, homology reveals the number of $k$-dimensional holes of a shape, for every integer $k$. For $k=0,1,2$, this corresponds to the number of connected components, tunnels, and cavities in the shape. Crucially, given a continuous map between two shapes (or, in this case, simplicial complexes), for instance, an inclusion from $X$ into $Y$, there is a well-defined map between the holes. Therefore, we can study the evolution of the homology in the filtration we obtained from data. The resulting evolution of topological features can be represented as a barcode (aka persistence diagram), a collection of intervals (bars) that represent the lifetime of a hole within the filtration. The length of a bar is called the persistence of the corresponding topological feature.
@@
A drawback of PH is that, unlike Mapper or UMAP, it does not easily scale to very big datasets. On the positive side, it is more intuitive than Mapper and UMAP because usually there is a natural choice of parameter to filter. Moreover, since it requires less parameter tuning, PH is very apt for automation.
@@ 4.3 Filtering results
After this filtering procedure, only the remaining publications are further considered for this survey. In total, over $7000$ results were screened, resulting in $35$ publications that were considered to be relevant for our stated research questions. These publications were then screened in detail. Based on this screening, the publications were categorized manually into different categories, that are presented in the following.
@@ 5 Results
In total $35$ works were identified as relevant for this survey. Each of those identified works was assigned to one of three clusters (A-C), based on its application within the production process. The identified clusters are:
@@
Cluster TDA Methods PH Mapper UMAP A  [64, 65, 66, 67, 42, 68, 69, 70, 71]  [72]  [73, 74, 75] B  [76, 77, 78, 41, 79, 80, 81]  [82, 83] C  [84, 85, 86, 87]  [47]  [52, 88, 89, 90, 91, 92, 93, 94]
@@
Cluster Work Input Data Format fin. $\subset\mathbb{R}^{n}$ $\mathbb{N}\to\mathbb{R}^{n}$ Scalar Field ($\mathbb{R}^{2}\to\mathbb{R}$) Graph Point Cloud Time Series Wafer Map Sign. Dist. Func. Microscope Imag. Surface Logs Funct. Block Netw. Task Graph A  [42] PH  [64] PH  [66] PH  [67] PH  [65] PH  [68] PH  [69] PH  [70] PH  [71] PH  [72] Mapper  [73] UMAP  [74] UMAP  [75] UMAP B  [41] PH  [76] PH  [77] PH  [78] PH  [79] PH  [82] Mapper  [83] Mapper  [80] PH  [81] PH C  [47] Mapper  [52] UMAP  [84] PH  [85] PH  [86] PH  [88] UMAP  [89] UMAP  [90] UMAP  [91] UMAP  [92] UMAP  [87] PH  [93] UMAP  [94] UMAP Count $7$ $17$ $3$ $1$ $2$ $2$ $1$ $1$ $1$ $8$ $2$
@@ 5.1 Quality Control on Product Level
For the segmentation of shapes in point cloud data, [42] propose a new method employing PH and graph convolutional networks. Their approach enriches the “local” features, extracted using the graph convolutional network, by “global” features. These features are extracted using PH, transformed to PIs and reshaped to a “topological feature map”. Their PH-based graph convolution network outperforms the state of the art on methods of fine-grained 3D shape segmentation, performed on point cloud data.
@@
Similarly, [73] analyzes the root cause in manufacturing variations of mechanical components during a production process. For this task, the input data is a 3D point cloud, acquired from an optical scanning step. With their work, they apply UMAP as a manifold learning method towards this task.
@@
On the task of additive manufacturing of radio-frequency devices, [75] propose a method for the identification of defect mechanism and their performance impact. This task usually requires in-line electromagnetic simulations, which are time-consuming and expensive. In their work, they use convolutional neural networks on microscope images, resulting in a vector. UMAP is then employed for the dimension reduction on the resulting vectors. Their approach provides to a faster and cheaper quality control process.
@@ 5.3 Manufacturing Engineering
A use-case on the planning of a multi-robot collective transport tasks is described in [87]. This work uses topological abstractional features for graph reinforcement learning. PH is used to derive higher-order topological information from an input task graph, to facilitate a feature vector for the policy network. The evaluation reports similar performance to state-of-the-art methods, while the necessary computational time was reduced significantly.
@@ 6.1 Interpretation of the Results
Method-wise, we observe a disparity in utilization. The most predominantly utilized method is PH (20 publications). This prevalence might be caused by its interpretability compared to the other methods. The dimension-reduction method UMAP is used in 11 referenced works. All of them underscore UMAPs topology-preserving characteristics, favoring its use over other dimension-reduction methods. The method used the least is Mapper with only 4 works. We speculate that a reason for this is that Mapper has been patented and commercialized [107], possibly preventing the academic community to freely study its applications. At the same time, the need to protect a TDA method proofs its commercial value.
@@
Unlike the others, Section 5.2 mostly (with one exception) leverages time series data. This is rather natural, since on process level is the level where time series data is most prevalent. However, the information about the process itself can be also embedded in a different format, e.g., as summary information of a production process, similar to what is being offered by the OPC UA specification [108] for Injection Moulding Machines to MES systems. In this case, the process information could also be embedded as a graph or a point cloud. This approach can be observed in the only, relatively recent work that is not dealing with time series data.
@@ 6.2 Opportunities and Challenges
UMAP helps against the curse of dimensionality, a prevalent challenge in data analysis characterized by the sparsity of data in high-dimensional spaces.
PH provides novel techniques to automate analytical processes by considering the full value range of a parameter. This approach is highlighted, for example, in the findings of a survey presented in [43].
UMAP and Mapper can deal with large datasets, a particularly valuable property in industrial settings, where data are usually big.
In general, TDA methods are more transparent than, for example, ML, where the analysis happens in a “black box”. Businesses typically prefer methods that are reliable over those that might offer better accuracy but are less understandable. From a strategic standpoint, this makes TDA a compelling choice for the application in industrial settings.
Moreover, TDA, by its very nature, extracts or uses information about shapes and topological features that no other data analysis method can find. However, one does not need to restrict to TDA, and it can be seamlessly integrated with other methodologies. An illustrative example is its combination with techniques fromML, where topological features are extracted using techniques from TDA. According to [21], the application of TDA with ML can be grouped into two main categories, namely intrinsic- and extrinsic topological features for ML. Intrinsic topological features incorporate topological analysis of the machine learning model itself and influence the model’s architecture or training. Interestingly, none of the identified works of this survey is of this category. We see this gap as potential opportunity for future TDA research in this application domain. Extrinsic topological features, however, enable the use of topological features extracted from a given dataset. This is enabled by vectorized persistence diagrams, like PI [54] or specialized kernels [56]. The degree of facilitating the topological information can vary and highly depends on the use-case and data:
@@
    Topological features are sufficient on its own, i.e., naive and direct analysis of vectorized representations of persistence diagrams. An example from this survey is the supervised analysis of chatter on time-series signals [78].
@@
    Adding topological information to existing pipelines. Information about shape is often informative and topological features can be used as complementary features to improve existing methods. The work in [87] illustrates this on a use-case by adding topological features to an already existing method and reporting an improved performance.
Another noteworthy approach to combine ML and TDA is, that ML methods preprocess data, so that topological aspects can be derived from the resulting representation. As an example, [72] facilitate a vision transformer to extract higher-level features from wafer maps, which are then analyzed using Mapper.
@@
    Time Series Data: For analyzes involving time series, PH is recommended due to its effectiveness in capturing temporal topological features. For dimension-reduction of multivariate time-series data, UMAP is advisable, as it preserves the topological structure of the data.
@@
    Point Cloud Data: If the point cloud is in low ambient dimension (up to 2) then PH, with the so-called alpha complex, is extremely efficient also for big datasets ($10^{4}$ points). In higher ambient dimension, PH is only feasible, so far, for datasets of $10^{2}$ or $10^{3}$ points, especially if one is interested in topological features of higher dimension (such as cavities), and not only low dimension (such as connected components and loops). For larger datasets, Mapper or UMAP should be considered first. The choice between them should be guided by the specific characteristics and requirements of the data.
@@
    Automation of Processes: For the automation of tasks or processes, anytime there is a clear choice for a varying parameter PH should be considered.
@@
    High-Dimensional Data: For data residing in high ambient dimensions, UMAP is the preferred method to mitigate the curse of dimensionality.
@@
We hope, moreover, that it will encourage practitioners in areas where PH, UMAP, or Mapper have yet to be used to test these methods. We anticipate that several tasks can be analyzed with TDA methods, resulting novel insights in these areas. Even problems considered “solved” (refer to [110] for examples and datasets) can be re-examined and further explored.

### v3

@@ Abstract
journal: Journal of Manufacturing Systems
@@ 3 Topological Data Analysis
We consider three main methods in TDA: Mapper [38], PH [17, 39], and Uniform Manifold Approximation and Projection (UMAP) [40]. PH is the prevalent method in TDA, Mapper a topologically guided clustering algorithm, and UMAP a topology-driven dimensionality reduction method. Common to all of them is that (i) the data at hand is converted into a suitable (geo)metric representation and (ii) topological properties are analyzed to extract (novel) information from the given representation. In a remote analogy, where the Fourier analysis tells us something about spectral properties of a signal, TDA extracts information about the shape of various kind of data (e.g., loops, tunnels, cavities, components, et cetera).
@@
During the stage of data preprocessing, data usually is manipulated by a set of parameters. The three aforementioned methods deal with parameters – i.e., the step from input data to a combinatorial object – in different ways. First, Mapper assembles parameters (and their values) in different groups and then clusters the inputs accordingly. This grouping can find latent relations in the dataset. Similarly, UMAP is a dimensionality reduction method. By projecting the data into lower-dimensional ambient space, redundant or not relevant information in the data is compressed. This enables the input to be analyzed more easily. On the other hand, PH overcomes the need of choosing parameter values by considering all possible values at the same time. Instead of outputting information about the shapes in the data for a single parameter value, PH keeps track of how the shapes evolve depending on the variation of parameter(s). PH is especially suitable for automated production since it removes the need to cherry-pick a threshold (see for example [43]).
@@ 3.1 Mapper
    An interval cover, cover for brevity, of the image of $f$ (i.e., a collection of possibly overlapping intervals whose union covers the image of $f$),
@@
We now describe the steps of Mapper (illustrated in Figure 5): first, the input data $X$ is reduced in dimension using the filter function $f$. This function projects the input points into a low-dimensional space. Then one chooses a cover $\mathcal{I}$ of the image of $f$ in the low-dimensional space. The idea is that the projection can facilitate the grouping of points by removing redundant information. This grouping, i.e. the cover, is then lifted back to the input points: one takes the preimage of each interval in the cover, i.e., the collection of all data $x\in X$ such that $f(x)\in\mathcal{I}$ for an interval $\mathcal{I}$ in the cover. In other terms, all data points that are mapped to the same interval in the cover are grouped together in the domain. Note that, since the intervals in the cover may overlap (and they usually do), a point $x\in X$ can belong to the preimages of several intervals. Each of these preimages is clustered according to the chosen clustering algorithm. Finally, each of these clusters is represented as a vertex, and two vertices have an edge between them if there is at least one $x\in X$ that belongs to both corresponding clusters.
@@
In practice, the major obstacles to the use of Mapper are the choices of the filter function $f$, cover, and clustering algorithm: the interpretability of the outcome entirely depends on these choices. While a few standard choices are known, a domain expert is usually required to get meaningful insights out of the Mapper pipeline. Indeed, the steps of the Mapper algorithm have very few mathematical constrains (the filter function can be anything, and the cover needs only to account for all points and consist of connected “pieces”). This in turn means that there are many degrees of freedom in applying this method, and that the correct choices are simply the ones that give better or more complete answers.
Nevertheless, Mapper is a powerful, versatile method that can reveal covert connectivity in datasets.
@@ 3.2 Uniform Manifold Approximation and Projection (UMAP)
The work by Hsu et al. [52] exemplifies the application of UMAP. They analyze failures in System Level Tests (SLTs) on integrated circuits. The need for robust interpretative methodologies is driven by the increasing transistor density and complexity in circuit designs, where the high-dimensional nature of the associated data poses a significant challenge. To address this, Hsu et al. [52] project high-dimensional data into a lower-dimensional space using UMAP. The obtained lower-dimensional representation is analyzed with a clustering algorithm, isolating clusters of systematically failing SLTs. These clusters are then leveraged to construct a decision tree, enabling precise identification of failures in individual components. The clustering algorithm is very efficient but sensitive to the curse of dimensionality. Therefore, applying it directly to the data would lower the performances.
@@ 3.3 Persistent Homology (PH)
There are several methods to turn $X_{r}$ into a simplicial complex. Some, such as alpha, Vietoris–Rips, and Čech complexes, rely on covering the data with increasingly bigger good covers, i.e., sets whose intersections have nice mathematical properties, like being contractible and open or closed, to apply a nerve theorem. This theoretical result, in turn, guarantees that the data and the resulting combinatorial object have the same topological properties, and thus the soundness of the analysis (see [53] for a survey of formulations of the nerve theorem and its applications in TDA). Another type of filtration, often used for building cubical complexes in images analysis, is the sublevelset filtration. Given the input $X$, one defines a function $f\colon X\to\mathbb{R}$, and then takes all the points in $X$ mapped under $f$ to a value smaller or equal the parameter $r$, for all $r\in\mathbb{R}$. Formally, $f^{-1}((-\infty,r])=\{x\in X\ |\ f(x)\leq r\}$. In the case of an image, this function is typically the average of the RGB values of a pixel. The choice of the filtration is sometime dictated by the input type, but in general there are several possibilities depending on the needs at hand. We will not further elaborate on details, but this choice may affect the interpretation and usually impacts the runtime considerably. Therefore, it is important to choose the more appropriate method for each application to ensure efficient computation [54].
@@
Now that we have a suitable combinatorial object, we need to extract information from it and we do so using homology. Homology is a fundamental concept from algebraic topology, allowing us to identify shapes that cannot be continuously deformed into each other. An extensive treatment is beyond our scope (see for example [17] for more details). Informally, homology reveals the number of $k$-dimensional holes of a shape, for every integer $k$. For $k=0,1,2$, this corresponds to the number of connected components, tunnels, and cavities in the shape. Crucially, given a continuous map between two shapes (or, in this case, simplicial complexes), for instance, an inclusion from $X$ into $Y$, there is a well-defined map between the holes. Therefore, we can study the evolution of the homology in the filtration we obtained from data. The resulting evolution of topological features can be represented as a barcode (aka persistence diagram), a collection of intervals (bars) that represent the lifetime of a hole within the filtration. The length of a bar is called the persistence of the corresponding topological feature. The filtration computed is stored as a matrix (called the boundary matrix), and the barcode is computed via a special (i.e., with only specific operations allowed) Gaussian elimination (see, for example [55] for more details on the barcode computation). Crucially, the barcodes generated by PH are stable. Therefore, we can define a meaningful distance on the data and on barcodes such that the distance between any two inputs bounds the distance between their corresponding barcodes [56]. Additionally, also the barcode’s vectorization is stable [57]. In practice, stability means that similar inputs result in similar outputs, ensuring that the extract information is robust against small perturbations of the input.
@@
A drawback of PH is that, unlike Mapper or UMAP, it does not easily scale to very big datasets since it boils down to Gaussian elimination on very big matrices. On the positive side, it is more intuitive than Mapper and UMAP because usually there is a natural choice of parameter to filter. Moreover, since it requires less parameter tuning, PH is very apt for automation.
@@ 4.3 Filtering results
After this filtering procedure, only the remaining publications are further considered for this survey. In total, over $7000$ results were screened, resulting in $34$ publications that were considered to be relevant for our stated research questions. These publications were then screened in detail. Based on this screening, the publications were categorized manually into different categories, that are presented in the following.
@@ 5 Results
In total $34$ works were identified as relevant for this survey. Each of those identified works was assigned to one of three clusters (A-C), based on its application within the production process. The identified clusters are:
@@
Cluster TDA Methods PH Mapper UMAP A  [67, 68, 69, 70, 42, 71, 72, 73, 74]  [75]  [76, 77] B  [78, 79, 80, 41, 81, 82, 83]  [84, 85] C  [86, 87, 88, 89]  [47]  [52, 90, 91, 92, 93, 94, 95, 96]
@@
Cluster Work Input Data Format fin. $\subset\mathbb{R}^{n}$ $\mathbb{N}\to\mathbb{R}^{n}$ Scalar Field ($\mathbb{R}^{2}\to\mathbb{R}$) Graph Point Cloud Time Series Wafer Map Sign. Dist. Func. Microscope Imag. Surface Logs Funct. Block Netw. Task Graph A  [42] PH  [67] PH  [69] PH  [70] PH  [68] PH  [71] PH  [72] PH  [73] PH  [74] PH  [75] Mapper  [76] UMAP  [77] UMAP B  [41] PH  [78] PH  [79] PH  [80] PH  [81] PH  [84] Mapper  [85] Mapper  [82] PH  [83] PH C  [47] Mapper  [52] UMAP  [86] PH  [87] PH  [88] PH  [90] UMAP  [91] UMAP  [92] UMAP  [93] UMAP  [94] UMAP  [89] PH  [95] UMAP  [96] UMAP Count $6$ $17$ $3$ $1$ $2$ $2$ $1$ $1$ $1$ $8$ $2$
@@ 5.1 Quality Control on Product Level
For the segmentation of shapes in point cloud data, [42] propose to enrich the input to an existing Graph Neural Network (GNN) with topological information. As GNNs typically do not capture multi-scale topological information, and misses these features in the input space. The authors propose an approach, that enriches the “local” features, extracted using a GNN (in fact, a Graph Convolutional Network), by “global” topological features. These features are extracted using PH, transformed to PIs and reshaped to a “topological feature map”. Both the local and the topological feature map are concatenated and fed to a $1\times 1$ convolutional filter to extract an “augmented feature map”. The augmented feature map is then used for the final segmentation task by the GNN. Their PH-based GNN outperforms the state of the art on methods of fine-grained 3D shape segmentation, performed on point cloud data.
@@
On the task of additive manufacturing of radio-frequency devices, [77] propose a method for the identification of defect mechanism and their performance impact. This task usually requires in-line electromagnetic simulations, which are time-consuming and expensive. In their work, they use Convolutional Neural Networks (CNNs) on microscope images, resulting in a vector. UMAP is then employed for the dimension reduction on the resulting vectors. Their approach provides to a faster and cheaper quality control process.
@@ 5.3 Manufacturing Engineering
A use-case on the planning of a multi-robot collective transport tasks is described in [89]. This work facilitates topological abstractional features for graph reinforcement learning. In their work, they replace the graph laplacian (as needed by their base method) with a laplacian matrix that is computed using PH. This laplacian matrix incorporates higher-order topological information from the input task graph, which is then used as an input for the policy network. The evaluation reports similar performance to state-of-the-art methods, while the necessary computational time was reduced significantly.
@@ 6.1 Interpretation of the Results
Method-wise, we observe a disparity in utilization. The most predominantly utilized method is PH (20 publications). This prevalence might be caused by its interpretability compared to the other methods. The dimension-reduction method UMAP is used in 10 referenced works. All of them underscore UMAPs topology-preserving characteristics, favoring its use over other dimension-reduction methods. The method used the least is Mapper with only 4 works. We speculate that a reason for this is that Mapper has been patented and commercialized [109], possibly preventing the academic community to freely study its applications. At the same time, the need to protect a TDA method proofs its commercial value.
@@
Unlike the others, Section 5.2 mostly (with one exception) leverages time series data. This is rather naturally caused by the manual clustering since the process level is the level where time series data is most prevalent. However, the information about the process itself can be also embedded in a different format, e.g., as summary information of a production process, similar to what is being offered by the OPC UA specification [110] for Injection Moulding Machines to MES systems. In this case, the process information could also be embedded as a graph or a point cloud. This approach can be observed in the only, relatively recent work that is not dealing with time series data.
In each of the three identified application cluster, the TDA methods utilize different characteristics of the input data. The works on Quality Control on Product Level (Section 5.1) are mainly on point cloud data. On this kind of data, PH is used to extract topological features and outperforms traditional methods, due to its ability to capture the data’s topological structure over multiple scales, while also being robust to noise (see Section 3.3). For cluster B, described in Section 5.2, mostly two characteristics of the data are leveraged: time series data are converted into a delayed embedding that reveals a hidden, yet different topological perspective , and Mapper is used to cluster the data according to its topological structure (see Sections 3.3 and 3.1, respectively). Further, the application of PH in the field on chatter detection reduces the influence of irrelevant noise. Within the cluster of Manufacturing Engineering (Section 5.3), the most used method is topological aware dimension reduction. Further, also in this case, Mapper reveals latent structure of high-dimensional process data (see Section 3.1). Additionally, PH is leveraged for its ability to automate processes by considering the full value range of a parameter and being robust to noise (see Section 3.3).
@@ 6.2 Opportunities and Challenges
UMAP helps against the curse of dimensionality, a prevalent challenge in data analysis characterized by the sparsity of data in high-dimensional spaces. Such kind of data can very likely be observed on production processes, where many sensory inputs are recorded (e.g. [110]).
PH provides novel techniques to automate analytical processes by considering the full value range of a parameter. This approach is highlighted, for example, in the findings of a survey presented in [43]. An example from the conducted survey is the detection of changes is a time series signal that highly depends on the value of the underlying signal, as with [26].
UMAP and Mapper can deal with large datasets, a particularly valuable property in industrial settings, where data are usually big (e.g. [110]).
In general, TDA methods are more transparent than, for example, ML, where the analysis happens in a “black box”. Businesses typically prefer methods that are reliable over those that might offer better accuracy but are less understandable. From a strategic standpoint, this makes TDA a compelling choice for the application in industrial settings, where explainability is often necessary for, e.g., quality control or certification purposes.
Opportunities by the Complementarity to other methods
Compared to other well established methods, TDA can overcome limitations in areas that are considered to be “solved”. One prime example are CNNs, where, depending on the fixed size of the kernel, only local information is extracted. Global information, like the shape of the captured object, is not directly accessible. Another noteworthy example are GNNs: Graphs and simplicial complexes share characteristics, as both can be considered as special cases of hyper-graphs [111, 112]. While GNNs have already shown significant success in various applications [113, 114], they are - by their design - inherently limited to pairwise interactions between nodes and fail to recognize higher-dimensional features, such as loops or cavities within graphs. With both examples, when leveraging a filtration procedure, as used in PH, this limitation can be overcome: by varying a parameter under observation over multiple scales, the topological information is extracted continuously from a local to global observation point of view. For a more detailed discussion on the relation of GNNs and TDA, we refer to [21, 115].
Moreover, TDA, by its very nature, extracts or uses information about shapes and topological features that no other data analysis method can find. Such an application is exemplified in [67], where changes in topological structures are detected in additive manufacturing. However, one does not need to restrict to TDA, and it can be seamlessly integrated with other methodologies. An illustrative example is its combination with techniques from ML (like Support Vector Machines, CNNs, or Transformers), where topological features are extracted using techniques from TDA. According to [21], the application of TDA with ML can be grouped into two main categories, namely intrinsic- and extrinsic topological features for ML. Intrinsic topological features incorporate topological analysis of the machine learning model itself and influence the model’s architecture or training. Interestingly, none of the identified works of this survey is of this category. We see this gap as potential opportunity for future TDA research in this application domain. Extrinsic topological features, however, enable the use of topological features extracted from a given dataset. This is enabled by vectorized persistence diagrams, like PI [58] or specialized kernels [60]. The degree of facilitating the topological information can vary and highly depends on the use-case and data:
@@
    Topological features are sufficient on its own, i.e., naive and direct analysis of vectorized representations of persistence diagrams. Here, no other features are used. An example from this survey is the supervised analysis of chatter on time-series signals [80].
@@
    Adding topological information to existing pipelines. Information about shape is often informative and topological features can be used as complementary features to improve existing methods. This can be accomplished by, e.g., concatenating topological features to the existing, classical input features. The work in [89] illustrates this on a use-case by adding topological features to an already existing method and reporting an improved performance.
Another noteworthy approach to combine ML and TDA is that ML methods preprocess data, so that topological aspects can be derived from the resulting representation. As an example, [75] facilitate a vision transformer to extract higher-level features from wafer maps, which are then analyzed using Mapper.
@@
    Time Series Data: For analyzes involving time series, such as industrial process data (see Section 5.2), PH is recommended due to its effectiveness in capturing temporal topological features [20]. For dimension-reduction of multivariate time-series data, UMAP is advisable, as it preserves the topological structure of the data [117].
@@
    Point Cloud Data: In higher ambient dimensions, the practical feasibility of computing PH with current implementations is typically limited to datasets containing $10^{2}$ to $10^{3}$ points (see the discussion on open challenges in TDA, Section 6). This limitation is particularly significant when analyzing higher-dimensional topological features, such as cavities, rather than just low-dimensional features like connected components and loops [62, 55, 54].
    If the point cloud is in low ambient dimension (up to 3; e.g., 2D images or 3D scans, like reported in Section 5.1), then PH, with the so-called alpha complex, is efficient also for bigger datasets ($10^{4}$ points) [118]. For larger datasets, Mapper or UMAP should be considered for dimension reduction first. The choice between them should be guided by the specific characteristics and requirements of the data.
@@
    Automation of Processes: For the automation of tasks or processes, anytime there is a clear choice for a varying parameter (e.g. analysis of signed distance functions or task graphs in Section 5.3) PH should be considered [119].
@@
    High-Dimensional Data: For data residing in high ambient dimensions, UMAP is the preferred method to mitigate the curse of dimensionality [49]. Sample applications can be found in Sections 5.1 and 5.3.
@@
We hope, moreover, that it will encourage practitioners in areas where PHUMAP, or Mapper have yet to be used to test these methods. We anticipate that several tasks can be analyzed with TDA methods, resulting novel insights in these areas. Even problems considered “solved” (refer to [120] for examples and datasets) can be re-examined and further explored.